### v2

@@ 3.5 Slider Crank
    A slider (body 3) with mass $m_{3}=1$ kg and moment of inertia $I_{3}=0.1\text{ kg}\cdot\text{m}^{2}$ is joined to the connecting rod by a revolute joint. It is also connected to a fixed wall via a spring with a torsional constant of $k=1$ N/m.
FNODE learns a mapping from the full 18-dimensional state space $(q,\dot{q})$ to accelerations $\ddot{q}$. Since the mechanism has only one DOF, the dynamics can be represented compactly with the minimal coordinates $(\theta_{1},\dot{\theta}_{1})$, with all other states dependent. The model hyperparameters are summarized in Table 5.
@@
Max.epochs 10000 400 5000 3000
@@
Figure 11 shows the slider-crank responses. All models fit the training data, but diverge in extrapolation. FNODE maintains accurate long-term predictions. MBD-NODE, while structurally similar to FNODE, is less stable and incurs higher computational cost due to its solver-based training. In contrast, LSTM and FCNN fail to generalize beyond the training window. Figure 15 illustrates the remaining state variables, which inherit errors from the minimal coordinates $(\theta_{1},\dot{\theta}_{1})$.
@@
Figure 11: Long-term prediction results for the slider-crank’s independent coordinate ($\theta_{1}$). The left column presents angular position versus time; the right column presents angular velocity versus time. Model performance on the training interval ($t\in[0,15]$) is shown with dashed lines, while extrapolated performance on the test interval is shown with dotted lines. The corresponding MSEs are: FNODE ($\epsilon=1.9\text{e-}3$), MBD-NODE ($\epsilon=8.2\text{e-}2$), LSTM ($\epsilon=4.5\text{e-}2$), and FCNN ($\epsilon=2.03\text{e}0$).
3.6 Cart Pole
[…]
Figure 16: Cart-pole system.
3.7 Cart-Pole System
We finally evaluate model performance on the cart-pole system, a canonical control benchmark (Fig. 16). The cart moves horizontally on a frictionless track, while the pendulum rotates freely about its pivot, with motion restricted to $\pi/2,\pi/2]$. The system state is described by four variables: cart position $x$, velocity $v$, pendulum angle $\theta$, and angular velocity $\omega$. The dynamics are governed by the nonlinear ODEs in Eqs. (27).
$x(0)=1,\;v(0)=0,\;\theta(0)=\tfrac{\pi}{6},\;\omega(0)=0.$
Table 6: Hyper-parameters for the cart-pole system.
@@ 3.7 Cart-Pole System
FNODE MBD-NODE LSTM FCNN
No. of hidden layers 3 3 3 3
No. of nodes per hidden layer 256 256 256 256
Max.epochs 10000 400 5000 3000
Initial learning rate 1e-3 1e-3 1e-3 1e-3
Learning rate decay 0.98 0.98 0.98 0.98
Activation function Tanh Tanh Tanh Tanh
Loss function MSE MSE MSE MSE
Optimizer Adam Adam Adam Adam
Figure 17: Phase-space trajectories for the cart-pole free-evolution scenario. The left subplot shows the cart’s phase portrait ($v$ vs. $x$), while the right subplot shows the pole’s phase portrait ($\omega$ vs. $\theta$). Dashed lines represent the fit to training data; dotted lines represent test data predictions. The models are: (a,b) FNODE, (c,d) MBD-NODE, (e,f) LSTM, and (g,h) FCNN.
Figure 18: State variable trajectories over time for the cart-pole system. The left column shows position variables ($x,\theta$), and the right column shows velocity variables ($v,\omega$). Dashed lines correspond to the training interval ($t\in[0,2]$), and dotted lines to the test interval. The final MSEs are: FNODE ($\epsilon=1.4\text{e-}5$), MBD-NODE ($\epsilon=8.6\text{e-}6$), LSTM ($\epsilon=6.6\text{e}0$), and FCNN ($\epsilon=1.4\text{e}0$).
Figures 17-18 compare model performance. FNODE and MBD-NODE both match the ground truth closely, with low MSEs ($1.4\!\times\!10^{-5}$ and $8.6\!\times\!10^{-6}$, respectively). MBD-NODE is slightly more accurate, but FNODE remains highly competitive while training far more efficiently (Table 7). In contrast, LSTM largely reproduces training patterns without generalization, and FCNN diverges significantly in extrapolation.
We also test FNODE in a model predictive control (MPC) task, where an external force $u$ is applied to stabilize the pole and return the cart to the origin. MPC proceeds by linearizing the dynamics at each step and solving a quadratic program over a finite horizon. For a linearized system $\dot{z}=Az+Bu$, the optimization problem is given in Eq. (28).
$\min_{u}$ $\sum_{k=0}^{N-1}z_{k}^{T}Qz_{k}+u_{k}^{T}Ru_{k}$ (28)
s.t. $z_{k+1}=Az_{k}+Bu_{k},\quad k=0,1,\dots,N-1$
$z_{k}\in Z,\quad u_{k}\in U,\quad k=0,1,\dots,N-1$
At each step, the system state is $z_{k}=(\theta_{k},x_{k},\omega_{k},v_{k})$ with control input $u_{k}$. The MPC problem is solved over a horizon of length $N$, subject to state and input constraints $\mathcal{Z}$ and $\mathcal{U}$. The cost function weights are $Q$ and $R$, both chosen as identity matrices. The linearized dynamics are written as $\dot{z}=Az+Bu$, where $A$ and $B$ are obtained from a first-order Taylor expansion of the nonlinear cart–pole equations (Eq. 29).
$A=\begin{bmatrix}0&0&1&0\\ 0&0&0&1\\ \frac{g(m+M)}{Ml}&0&0&0\\ -\frac{mg}{M}&0&0&0\end{bmatrix},\quad B=\begin{bmatrix}0\\ 0\\ -\frac{1}{Ml}\\ \frac{1}{M}\end{bmatrix}$ (29)
A trained FNODE model can be directly linearized by computing the Jacobian of its learned dynamics. In this task, FNODE learns the mapping from state and control input to accelerations, $f(\theta_{k},x_{k},\omega_{k},v_{k},u)\mapsto(\ddot{\theta}_{k},\ddot{x}_{k})$. The Jacobian is obtained via automatic differentiation and used as the linearized matrices $A$ and $B$ within the MPC framework.
Figure 19 shows the MPC results for the cart-pole. The five subplots depict the pole angle $\theta$, cart position $x$, angular velocity $\omega$, cart velocity $\dot{x}$, and control input $u$. The FNODE-based controller (red dotted) closely matches the analytical baseline (blue solid), demonstrating that FNODE learns the dynamics with sufficient accuracy for MPC.
Figure 19: Comparative performance of the MPC strategy. The analytical model (solid blue) and FNODE-based model (dotted red) are compared across: (a) pole angle $\theta$, (b) cart position $x$, (c) pole angular velocity $\omega$, (d) cart velocity $v$, and (e) the control force $u$.
@@ 4 Discussion
Tables 7 and 8 summarize the training cost and predictive accuracy of the four models across all benchmarks. Two main observations emerge.
First, FNODE consistently trains much faster than MBD-NODE, often by one to two orders of magnitude. This gap arises because FNODE bypasses the adjoint sensitivity method and does not embed an ODE solver in the backpropagation loop. On larger problems such as the slider-crank, the efficiency gain is especially pronounced: FNODE completes training in $\sim$119 s versus over 2700 s for MBD-NODE. Compared to purely data-driven baselines (FCNN, LSTM), FNODE’s training times are of the same order of magnitude, demonstrating that its physics-informed formulation does not incur significant overhead.
[…]
Ultimately, FNODE provides a favorable balance between efficiency and accuracy: it is orders of magnitude faster to train than MBD-NODE while delivering consistently higher fidelity than LSTM and FCNN, and near-parity or better accuracy than MBD-NODE across benchmarks. These results highlight FNODE’s potential as a scalable framework for data-driven simulation of constrained multibody systems.
@@
Single Mass Spring Damper FNODE 18.17
MBD-NODE RK4 540.18
FCNN 16.96
LSTM 500.63
Triple Mass Spring Damper FNODE 11.45
MBD-NODE RK4 529.86
FCNN 9.67
LSTM 5.73
Double Pendulum FNODE 15.04
MBD-NODE RK4 553.96
FCNN 8.94
LSTM 5.66
Slider Crank FNODE 118.98
MBD-NODE RK4 2735.80
FCNN 27.81
LSTM 33.05
Cart Pole FNODE 14.70
MBD-NODE RK4 167.87
FCNN 6.52
LSTM 6.98
@@
Single Mass-Spring-Damper 1.9e-2 1.3e-1 7.0e-1 4.2e1
Triple Mass-Spring-Damper 3.0e-2 5.0e-2 1.0e0 1.3e0
Double Pendulum 1.9e-2 3.7e-1 2.2e0 8.7e0
Slider Crank 1.9e-3 8.2e-2 4.5e-2 2.03e0
Cart Pole 1.4e-5 8.6e-6 6.6e0 1.4e0

### v3

@@ 3.5 Slider Crank with Coulomb friction
    A slider (body 3) with mass $m_{3}=1$ kg and moment of inertia $I_{3}=0.1\text{ kg}\cdot\text{m}^{2}$ is joined to the connecting rod by a revolute joint. It is also connected to a fixed wall via a spring with a torsional constant of $k=1$ Nm/rad. There is friction between the slider and the ground, which is modeled as Coulomb friction with a coefficient of $\mu$.
FNODE learns a mapping from the full 18-dimensional state space $(q,\dot{q})$ to accelerations $\ddot{q}$. Since the mechanism has only one DOF, the dynamics can be represented compactly with the minimal coordinates $(\theta_{1},\dot{\theta}_{1})$, with all other states dependent. Following the coordinate-partitioning approach [43], once FNODE produces the acceleration of the crank, numerical integration is used to compute the orientation and the angular velocity of the crank, while the kinematic constraint equations are used to recover the dependent generalized coordinates. The model hyperparameters are summarized in Table 6.
@@
Max. epochs 450 300 300 450
Initial learning rate 5e-4 1e-3 1e-3 1e-3
Learning rate decay 0.99 0.98 0.98 0.98
Activation function tanh tanh Sigmoid,tanh tanh
Loss function MSE MSE MSE MSE
Optimizer Adam Adam Adam Adam
3.5.1 Non-friction Case
Figure 12 shows the slider-crank responses. All models fit the training data, but diverge in extrapolation. FNODE maintains accurate long-term predictions. MBD-NODE, while structurally similar to FNODE, is less stable and incurs higher computational cost due to its solver-based training. In contrast, LSTM and FCNN fail to generalize beyond the training window. Figures 1316 illustrate the remaining state variables, which inherit errors from the minimal coordinates $(\theta_{1},\dot{\theta}_{1})$.
Figure 12: Long-term prediction results for the frictionless slider-crank’s independent coordinate ($\theta_{1}$). The left column presents angular position versus time; the right column presents angular velocity versus time. Model performance on the training interval ($t\in[0,15]$) is shown with dashed lines, while extrapolated performance on the test interval is shown with dotted lines. The corresponding MSEs are: FNODE ($\epsilon=1.7\text{e-}1$), MBD-NODE ($\epsilon=8.7\text{e-}1$), LSTM ($\epsilon=1.4\text{e}0$), and FCNN ($\epsilon=3.1\text{e}0$).
3.5.2 Friction Case
In this experiment, the friction coefficient $\mu$ is added as an input to the FNODE model to test the interpolation in the parameter space. The friction coefficient is sampled from $[0.0,0.6]$ uniformly for training data generation. For testing, we use interpolated friction coefficients to test the generalization ability of FNODE in parameter space.
Figure 17 shows FNODE-predicted phase-space trajectories of the slider-crank mechanism for different friction coefficients. The FNODE model was trained on coefficients sampled uniformly from $[0.0,0.6]$ and evaluated on interpolated values. This test case indicates that FNODE can learn the frictional dynamics and approximate the resulting non-smooth acceleration. To improve fidelity near these non-smooth transitions, one can increase FNODE model complexity (e.g., deeper/wider hidden layers) to better represent piecewise-smooth behavior.
Figure 17: Trajectories of the slider-crank mechanism under varying friction coefficients, as predicted by the FNODE model. The model was trained on friction coefficients sampled uniformly from $[0.0,0.6]$ and tested on interpolated values. In this figure, c = 0.0 and c = 0.6 are from the training set, while c = 0.15, 0.35 are interpolated test cases.
A central challenge for FNODE modeling in the presence of friction is accurately predicting the resulting non-smooth acceleration field. A common way to improve accuracy in such settings is to increase network depth, which enhances the model’s ability to represent sharp, piecewise-smooth transitions. Figure 18 shows FNODE acceleration approximations for the slider-crank mechanism at $\mu=0.6$ using models with different numbers of hidden layers. As depth increases, the approximation consistently improves, as evidenced by the decreasing MSE values reported in Table 7.
Figure 18: Acceleration prediction for the slider-crank mechanism ($\mu=0.6$) using FNODE models with varying depths.
Table 7: MSE of various-layers FNODE acceleration prediction for the slider-crank mechanism with friction coefficient $\mu=0.6$
Number of Hidden Layers MSE
2 3.70e-02
3 1.72e-02
4 1.47e-02
5 1.44e-02
6 1.41e-02
3.6 Cart-Pole System
Figure 19: Cart-pole system.
We conclude with the cart-pole system, considered here because it is a canonical control benchmark (Figure 19). The cart can move horizontally on a frictionless track, while the pendulum swings freely about its pivot, with motion restricted to $\pi/2,\pi/2]$. The system state is described by four variables: cart position $x$, velocity $\dot{x}$, pendulum angle $\theta$, and angular velocity $\dot{\theta}$. The dynamics are governed by nonlinear ODEs:
$x(0)=1,\;\dot{x}(0)=0,\;\theta(0)=\tfrac{\pi}{6},\;\dot{\theta}(0)=0.$
Table 8: Hyper-parameters for the cart-pole system.
Hyper-parameters Model
FNODE MBD-NODE LSTM FCNN
No. of hidden layers 2 2 2 2
No. of nodes per hidden layer 256 256 256 256
Max. epochs 450 400 300 450
@@ 3.6 Cart-Pole System
Figure 20: Phase-space trajectories for the cart-pole free-evolution scenario. The left subplot shows the cart’s phase portrait ($\dot{x}$ vs. $x$), while the right subplot shows the pole’s phase portrait ($\dot{\theta}$ vs. $\theta$). Dashed lines represent the fit to training data; dotted lines represent test data predictions. The models are: (a,b) FNODE, (c,d) MBD-NODE, (e,f) LSTM, and (g,h) FCNN.
@@
Figure 21: State variable trajectories over time for the cart-pole system. The left column shows position variables ($x,\theta$), and the right column shows velocity variables ($\dot{x},\dot{\theta}$). Dashed lines correspond to the training interval ($t\in[0,2]$), and dotted lines to the test interval. The final MSEs are: FNODE ($\epsilon=6.7\text{e-}4$), MBD-NODE ($\epsilon=1.1\text{e-}3$), LSTM ($\epsilon=1.6\text{e-}1$), and FCNN ($\epsilon=3.5\text{e-}1$).
Figures 20-21 compare model performance. FNODE and MBD-NODE both match the ground truth closely, with low MSEs ($6.7\!\times\!10^{-4}$ and $1.1\!\times\!10^{-3}$, respectively). MBD-NODE is slightly more accurate, but FNODE remains competitive while training far more efficiently. LSTM largely reproduces training patterns, which fit the periodic pattern of cart-pole system, so it still provides relatively accurate predictions. FCNN diverges significantly in extrapolation.
We also test FNODE in a model predictive control (MPC) task, where an external force $u$ is applied to stabilize the pole and return the cart to the origin. The MPC strategy, for control purposes, linearizes the dynamics and then solves a quadratic program over a finite horizon [52]. For a linearized system $\dot{z}=Az+Bu$, the optimization problem is formulated as:
$\min_{u}$ $\sum_{k=0}^{N-1}z_{k}^{T}Qz_{k}+u_{k}^{T}Ru_{k}$ (32)
s.t. $z_{k+1}=(I+\Delta tA)z_{k}+\Delta tBu_{k},\quad k=0,1,\dots,N-1$
$z_{k}\in\mathcal{Z},\quad u_{k}\in\mathcal{U},\quad k=0,1,\dots,N-1$
At each step, the system state is $z_{k}=(\theta_{k},x_{k},\dot{\theta}_{k},\dot{x}_{k})^{T}$ with control input $u_{k}$. The MPC problem is solved over a horizon of length $N$, subject to state and input constraints $\mathcal{Z}$ and $\mathcal{U}$. The cost function weights are $Q$ and $R$, both chosen as identity matrices. $A$ and $B$ are the linearized system matrices, obtained from the Jacobian of the dynamics.
To solve the control problem, FNODE learns the mapping from state and control input to accelerations, $f(\theta_{k},x_{k},\dot{x}_{k},\dot{\theta}_{k},u)\mapsto(\ddot{\theta}_{k},\ddot{x}_{k})$. We use automatic differentiation and the inherent differentiable nature of a Neural Network to obtain the linearization matrices $A$ and $B$ required by the MPC framework. The standard (non-NN based) MPC for Cart-Pole is also provided as a baseline for comparison, see D for the expression of the linearization matrices.
Figure 22 shows the MPC results for the cart-pole. The five subplots depict the pole angle $\theta$, cart position $x$, angular velocity $\dot{\theta}$, cart velocity $\dot{x}$, and control input $u$. The FNODE-based controller (red dotted) closely matches the analytical baseline (blue solid), demonstrating that FNODE learns the dynamics with sufficient accuracy for supporting an MPC controller.
Figure 22: Comparative performance of the MPC strategy. The analytical model (solid blue) and FNODE-based model (dotted red) are compared across: (a) pole angle $\theta$, (b) cart position $x$, (c) pole angular velocity $\dot{\theta}$, (d) cart velocity $\dot{x}$, and (e) the control force $u$.
3.7 Parameterized Vehicle Model
We finally consider the simplified 4DOF vehicle model proposed in [53, 54]. The model has a simple powertrain component and adopts throttle and steering as control inputs. The state and control vectors are:
$q=[x,\,y,\,\theta,\,v]^{T},\qquad u=[\alpha,\,\delta]^{T},$ (33)
where $(x,y)$ is the vehicle position in the global frame, $\theta$ is the heading (yaw) angle, $v$ is the longitudinal speed, $\alpha$ is the normalized throttle, and $\delta$ is the steering angle. The state transition model is
$\dot{q}=\begin{bmatrix}\dot{x}\\ \dot{y}\\ \dot{\theta}\\ \dot{v}\end{bmatrix}=\begin{bmatrix}\cos\theta\cdot v\\ \sin\theta\cdot v\\ \dfrac{v\tan\beta\delta}{l}\\[4.0pt] \dfrac{2\,T(\alpha,v)\,\gamma}{I_{\mathrm{wheel}}}\,R_{\mathrm{wheel}}\end{bmatrix}$, (34)
where $l$ is the wheelbase, $\beta$ maps steering command $\delta$ to a physical wheel steering angle, $\gamma$ is the gear ratio, and $I_{\mathrm{wheel}}$ and $R_{\mathrm{wheel}}$ are the wheel inertia and radius. The motor torque is modeled as
$T(\alpha,v)$ $=T^{\prime}(\alpha,\omega_{m})=\alpha f_{1}(\omega_{m})-c_{1}\omega_{m}-c_{0}\;,$ (35)
$f_{1}(\omega_{m})$ $=-\frac{\tau_{0}}{\omega_{0}}\omega_{m}+\tau_{0}\;,$
$\omega_{m}$ $=\frac{v}{R_{\mathrm{wheel}}\gamma}\;,$
where $\tau_{0}$ and $\omega_{0}$ are the stalling torque and no-load speed of the motor, and $c_{0},c_{1}$ model the motor resistance torque. This example contains nine scalar parameters
$\Theta_{p}=[l,\,\beta,\,\gamma,\,I_{\mathrm{wheel}},\,R_{\mathrm{wheel}},\,c_{0},\,c_{1},\,\tau_{0},\,\omega_{0}]^{T}\;,$ (36)
each sampled on a grid with four values, yielding $4^{9}$ parameter combinations. We simulate each parameter instance for 20 s with time step $\Delta t=0.01$ (2000 steps). The control signal $u(t)=[\alpha(t),\delta(t)]^{T}$ is pre-defined and shared by all parameter instances; the corresponding state trajectories are generated by the RK4 integrator. For each trajectory, the first 1500 steps are used for training and the remaining 500 steps for testing (extrapolation).
Table 9: Parameter ranges for the 4DOF vehicle model
Parameter Range Explanation
$l$ [2.7, 2.9] wheelbase
$\beta$ [0.60, 0.75] steering ratio
$\gamma$ [0.15, 0.25] gear ratio
$I_{\mathrm{wheel}}$ [0.9, 1.3] wheel inertia
$R_{\mathrm{wheel}}$ [0.32, 0.35] wheel radius
$c_{0}$ [0.015, 0.025] motor resistance torque coefficient
$c_{1}$ [0.025, 0.045] motor resistance torque coefficient
$\tau_{0}$ [250, 350] stalling torque
$\omega_{0}$ [1200, 1500] no-load speed
Table 10: Test parameters for the four test cases.
Test $l$ $R_{\mathrm{wheel}}$ $I_{\mathrm{wheel}}$ $\tau_{0}$ $\omega_{0}$ $c_{0}$ $c_{1}$ $\beta$ $\gamma$
1 2.7 0.32 0.9 250 1200 0.015 0.025 0.6 0.15
2 2.9 0.35 1.3 350 1500 0.025 0.045 0.75 0.25
3 2.726 0.3281 0.952 277 1239 0.0177 0.0276 0.6405 0.163
4 2.846 0.3461 1.192 337 1419 0.0237 0.0396 0.7305 0.223
FNODE learns the mapping from the full state and control inputs to the state time derivative, i.e., $(q,u,\Theta_{p})\mapsto\dot{q}$ as defined in Eqs. (34)–(36). All models are trained on the same trajectories and evaluated on the held-out extrapolation interval. The hyperparameters are summarized in Table 11.
Table 11: Hyper-parameters for the 4DOF vehicle model.
@@ 3.7 Parameterized Vehicle Model
FNODE
No. of hidden layers 6
No. of nodes per hidden layer 256
Max. epochs 300
Initial learning rate 1e-4
Learning rate decay 0.98
Activation function tanh
Loss function MSE
Optimizer AdamW
Figure 23: Trajectory Comparison of the 4DOF vehicle model.
Figure 23 compares the predicted trajectories of $q$ against the ground truth under identical control inputs. Tests 1-2 are drawn from the corners of the parameter grid, whereas tests 3-4 lie in the interior (Table 10) and therefore assess interpolation in parameter space. All methods track well within the training interval; discrepancies arise under extrapolation, where integration errors accumulate and parameter variations compound. FNODE remains stable across the grid and delivers comparatively accurate predictions over the test window, while the baselines exhibit larger drift and occasional divergence. Because the system’s high dimensionality makes the baseline datasets prohibitively large, we report only FNODE results here. Future work will investigate scalable baseline implementations for high-dimensional parameter spaces.
@@ 4 Discussion
Figure 24 summarizes the computational cost and predictive accuracy of the four models for the triple mass-spring-damper benchmark. Two observations emerge.
First, the training process of FNODE converges with lower computational cost than MBD-NODE. The computational cost is quantified by the training time of the model. As shown in Figure 24, the train loss of FNODE converges with lower training time compared to MBD-NODE. This gap arises because FNODE bypasses the adjoint sensitivity method and does not embed an ODE solver in the backpropagation loop. Compared to purely data-driven baselines such as FCNN, FNODE’s training times are of the same order of magnitude.
Second, FNODE achieves consistently lower prediction error than the black-box baselines and is competitive with, or superior to, MBD-NODE. In dissipative and chaotic systems (single/triple mass-spring-damper, double pendulum), FNODE outperforms all baselines by large margins, maintaining stable long-term predictions. For the slider-crank mechanism, FNODE achieves the best accuracy ($1.8\!\times\!10^{-3}$ MSE), significantly outperforming both MBD-NODE and the data-driven models. Only in the cart-pole case does MBD-NODE obtain a slightly lower error than FNODE. These results underscore the importance of accurate acceleration targets: FFT provides precision for smooth trajectories, while FD ensures robustness for non-periodic signals and trajectory endpoints.
FNODE provides a good efficiency vs. accuracy balance: it is orders of magnitude faster to train than MBD-NODE while delivering consistently higher fidelity than LSTM and FCNN, and near-parity or better accuracy than MBD-NODE across benchmarks. These results suggest that FNODE is computationally practical for higher-dimensional constrained multibody systems and can accommodate parameter variation, as illustrated by the added higher-DOF vehicle example and parameter-conditioned experiments.
In its current form, FNODE is best suited to systems with smooth or piecewise-smooth trajectories, where accelerations are well-defined and can be estimated reliably from sampled data (here via FFT/FD differentiation). Our results indicate that, despite the resulting non-smoothness, a Coulomb friction model remains amenable to FNODE, which can learn an accurate approximation of the associated acceleration field. In contrast, genuinely nonsmooth events, e.g., impacts (chattering) and complementarity-based stick-slip transitions, are deferred to future work, as discussed in the Conclusions.
@@
Single Mass Spring Damper FNODE 112.46
MBD-NODE RK4 201.90
FCNN 103.28
LSTM 171.99
Triple Mass Spring Damper FNODE 108.23
MBD-NODE RK4 244.80
FCNN 102.78
LSTM 172.05
Double Pendulum FNODE 89.33
MBD-NODE RK4 212.09
FCNN 88.04
LSTM 148.41
Slider Crank FNODE 565.01
MBD-NODE RK4 1185.64
FCNN 453.50
LSTM 813.11
Cart Pole FNODE 59.46
MBD-NODE RK4 144.51
FCNN 59.36
LSTM 96.08
4dof Vehicle FNODE 3304.32
MBD-NODE RK4
FCNN
LSTM
@@
Figure 24: The test MSE vs training time comparison of the all models on the triple mass-spring-damper benchmark.
@@
Single Mass-Spring-Damper 7.6e-5 1.4e-3 1.5e-2 1.6e-1
Triple Mass-Spring-Damper 6.1e-3 6.8e-3 2.5e-1 9.0e-2
Double Pendulum 1.4e-1 2.3e-1 7.6e-1 1.9e0
Slider Crank 1.7e-1 8.7e-1 1.4e0 3.1e0
Cart Pole 6.7e-4 1.1e-3 1.6e-1 3.5e-1